OpenStudy (anonymous):

need help solveing for n n x 3/5=1

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }n = 1\] multiply by the reciprocal to get n by itself and you'll have the answer

OpenStudy (anonymous):

so multipy by one

OpenStudy (anonymous):

multiply by the recipicol.

OpenStudy (anonymous):

No. Multiply 1 by the reciprocal of 3/5 which is 5/3

OpenStudy (anonymous):

To make this simple, the answer is going to be n = 1.67 (3/5)N = 1 N = (5/3)1 N = (5/3) which is also 1.67